03 March 2014

Cloudy day

Today was cloudy with light snow at Eureka. Temperatures outside remained moderate: Eureka was -26°C, while PEARL was -23°C. The Polar Vortex continues to be highly distorted and elongated from Quebec to Kamchatka, but is no longer directly over Eureka.

In the PEARL Ridge Lab's IR lab, Joseph and Dan were unable to take measurements with the Bruker or PARIS due to the cloudy weather. It was a quiet day in the IR lab, and so they worked on other research tasks.

Over in the UV-Vis Lab, Sophie analyzed the results of yesterday's tests on the UT-GBS. They showed an acceptable CCD alignment, so she performed stray light tests and polarization tests. Everything is looking good, so Sophie and Paul moved the UT-GBS' telescope into the hatch. The UT-GBS is measurement-ready! Sophie began measurements with the UT-GBS this afternoon. Dark current measurements will be recorded overnight. Depending on these results, she might fine-tune some parameters.

Today Tom and Zahra put the suntracker for SPS and MAESTRO into sun-tracking mode to correct for elevation and azimuth alignment. In order to correct for elevation and azimuth they had to manually rotate the tracker until it was looking at the Sun. They then made a sighting to ensure the incoming beam of light aligns with the internal aperture. In the process they discovered that the tracker winds up the cables and gets stuck. Tom and Zahra will have to remove the lid of the tracker and check if the tracker's reference point is pointing properly. They plan to do that tomorrow.

Tom and Zahra also moved Brewer #192 to the roof today and tried to test it. The computer had a problem speaking to the Brewer. They swapped its cables and COM ports. This did not help. They then restarted the tracker and Brewer. When restarted the tracker must find its zero and it starts rotating in order to do so. After restarting it failed to do this. They believe the firmware might have a problem. Tom and Zahra will reload it tomorrow and see if it fixes the problem.

Down at 0PAL, Emily and Sham ran the lidar overnight in all channels. Eventually, they had to close the hatch due to snow on the roof window. They took advantage of the hatch-closed time to make some dark count profiles in various configurations. Emily and Sham also built a support for a large polarizer, which they later installed over the primary mirror. Before supper, they tested it. After supper, they spent six hours doing a depolarization calibration. Without both of them at the CRL it would no doubt have taken three times as long! Zahra also visited the lab to make some burn paper "postcards" while learning about the CRL. Before heading off to bed, Emily and Sham set the lidar up to take regular measurements overnight, so long as the weather holds out.

Today's ACE ozonesonde flight occurred at 6:15 PM (local time) and reached an altitude of 31 438 meters (8.8 hPa).

U of T News has published two articles (here and here) about the campaign so far. If you want to read even more about what we do at PEARL, check it out and share!

The team is hoping for sunshine tomorrow.

- Dan
Back to Daily Reports
Back to Home